OpenStudy (anonymous):

what is the simplified form of... 1/10xy • 5x^2/7y

OpenStudy (callisto):

\[\frac{1}{10xy} • \frac{5x^2}{7y}\]\[=\frac{5x^2}{(7)10xy^2} \]\[=\frac{x}{(7)2y^2} \]\[=\frac{x}{14y^2}\]